OpenStudy (kpop4life123):

What is the volume of a cylinder with a radius of 7 feet and a height of 12 feet? Use 3.14 for pi. Round your answer to the nearest hundredth.

pooja195 (pooja195):

\[\huge~\rm~\bf V=\pi~r^2h\]

pooja195 (pooja195):

\[\huge~\rm~\bf V=(3.14)~(7)^ 2(12)\]

OpenStudy (anonymous):

It's the area of a circle multiplied the cylinder's height. \[\pi 7^2 12=1847.26 \]

OpenStudy (anonymous):

1846.32 using 3.14 for the value of pi.